OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. m =fraction 3/5 , P(3, 2) and Q(8,_)

OpenStudy (anonymous):

So they tell you your slope is = 3/5, right? So you'll need to set up an equation to find the missing y coordinate. Do you remember the slope formula I showed you?

OpenStudy (anonymous):

yea y-y x-x

OpenStudy (anonymous):

Okay, so they told you Y-Y/X-X = 3/5

OpenStudy (anonymous):

Go ahead and fill in the Y's and the X's

OpenStudy (anonymous):

x 8-3 y 2-?

OpenStudy (anonymous):

right, you have (Y-2)/(8-3) which is = (Y-2)/5. Now set (Y-2)/5 = 3/5

OpenStudy (anonymous):

How would you solve that equation?

OpenStudy (anonymous):

by dividing

OpenStudy (anonymous):

You have a 5 in the denominator. How would you get rid of that?

OpenStudy (anonymous):

subtracting?

OpenStudy (anonymous):

No. If you multiplied both sides by 5 you would have 5/5, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5(Y-2)/5 = 5*3/5

OpenStudy (anonymous):

So whats 5/5 equal to?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Right. So what are we left with if the 5's cancel?

OpenStudy (anonymous):

There is Y-2 on the left side. What about the right side?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

No. You have 1*(Y-2) = 1*3

OpenStudy (anonymous):

So you have Y-2 = ?

OpenStudy (anonymous):

What is 1 times 3? Or one times any number?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so you end up with Y-2=3. Can you solve that?

OpenStudy (anonymous):

so its 5-2=3

OpenStudy (anonymous):

that's correct. Y = 5

OpenStudy (anonymous):

thank you